OpenStudy (anonymous):

Write a rule for the linear function whose graph has slope m and contains the given point. m = -1/4; (4, 0)

OpenStudy (anwaarullah):

do you knoe slope formula??

OpenStudy (anwaarullah):

you can do it lol

OpenStudy (ybarrap):

 $$ m=slope\\ b=y-intercept $$ You can find \(b\) by setting x=4 and y=0: $$ 0=m\times 4+b $$ You know m, that's the slope (i.e m=-1/4). Now that you know \(b\), you can just plugin this in along with the slope into the following and you're done! $$ y=mx+b $$